Calculate the difference between the simple interest and the compound interest if R3 400 is invested at 10,5% p.a. for 6 years
1
Expert's answer
2020-04-08T16:55:04-0400

Simple interest:

"S=P(1+i\times n)=3400(1+0.105\times6)=5542"

Compound interest

"S=P(1+i)^n=3400(1+0.105)^6=6 189.36"

"\Delta=6189.36-5542=" 647.36